### v1

# Wigner function method for the Gibbons–Hawking and the Unruh effect

Abstract
An observer at rest with the expanding universe experiences some extra noise in the quantum vacuum, and so does an accelerated observer in a vacuum at rest (in Minkowski space). The literature mainly focuses on the ideal cases of exponential expansion (de–Sitter space) or uniform acceleration (Rindler trajectories) or both, but the real cosmic expansion is non–exponential and real accelerations are non–uniform. Here we use the frequency–time Wigner function of vacuum correlations to define time–dependent spectra. We found excellent Planck spectra for a class of realistic cosmological models, but also strongly non–Planckian, negative Wigner functions for a standard scenario testable with laboratory analogues.
Keywords: quantum vacuum, cosmic expansion, accelerated observers, laboratory analogues, time-dependent spectra.
1 Introduction
Imagine an observer at rest with the empty space of the expanding universe (Fig. 1). Suppose the observer is equipped with a sensitive spectrometer for detecting electromagnetic radiation. The spectrometer picks up the noise of vacuum fluctuations and also some extra noise that closely resembles the Hawking radiation of black holes [1, 2]. The extra noise depends on the expansion, and it is due to a difference in the measure of time the observer and the vacuum perceive. The spectrometer decomposes the field into frequency components with respect to the proper time of the observer, whereas the fluctuations of the field (Fig. 1) are continuously red–shifted by cosmic expansion [3]. A similar difference in time appears in the opposite scenario: an accelerated observer in a vacuum at rest [4, 5, 6] (in non–expanding Minkowski space). There the proper time of the observer differs from the time the vacuum experiences, too. The extra radiation is related to the Hawking radiation of black holes [1], because in both cases the observer is shrouded by horizons. In the expanding universe, it is the cosmological horizon [3] where the expansion velocity reaches the speed of light. In the case of the accelerated observer, it is the Rindler horizon [1] where the observer approaches the speed of light. For an exponentially expanding, spatially flat universe (de–Sitter space) the cosmological horizon is light–like [2] as the event horizon of an ideal black hole [1] and the observer would measure a perfect Planck spectrum [2]. But the real evolution of the universe is not exponential yet [7]. What is the spectrum there? The same question applies to the accelerated observer. A uniformly accelerated observer would measure a true Planck spectrum. But real accelerations must begin and end some time or may not be uniform at all. Is then the spectrum still Planckian?
Renaud Parentani [8, 9, 10] had worked out the precise condition when real black holes radiate a Planck spectrum. These real black holes are black–hole analogues in Bose–Einstein condensates [11]; they are real, because they are made in real experiments [11] and they correspond to a wide class of regimes [12] likely to affect the emission of Hawking radiation from astrophysical black holes. In an ideal gravitational black hole, the speed of light does not depend on wavelength, even at the event horizon where light would stand still, oscillating with wavelengths shorter than any scale. For real astrophysical black holes, the wave velocity will eventually change with wavelength [12], releasing light from the grip of the horizon. In Bose–Einstein condensates, the analogue of the speed light — the speed of sound — definitely depends on the wavelength [13]. Analogues of cosmic expansion have been made as well and in those analogues particle creation was seen [14, 15]. Almost arbitrary expansion histories of the universe can be tested [16, 17] with Bose–Einstein condensates [14]. The related case of radiation perceived by non–uniform acceleration has been a notoriously difficult theoretical problem [18, 19, 20, 21, 22, 23] but can be measured in analogues as well [24, 25].
In this paper, we hope to shed some light on the radiation spectrum in the expanding universe, known as the Gibbons–Hawking effect [2], for non–exponential expansion and on the radiation perceived by accelerated observers, the Fulling–Davies–Unruh effect [4, 5, 6] or Unruh effect in short, for non–uniform acceleration. The paper builds upon and corrects previous work [23, 26, 27] using the Wigner function [28, 29, 30] for defining and calculating time–dependent spectra. We found excellent agreement with a Planck spectrum for one class of models, and poor agreement for another. What the condition for thermality and the effective temperatures are remains a mystery; it would probably take someone like Renaud Parentani to clarify it.
2 Correlation function
Consider a spatially flat universe expanding with scale factor $a(t)$ where $t$ denotes the time perceived by an observer at rest, the cosmological time. The quantum vacuum propagates with conformal time [7]:
$\tau=\int\frac{\mathrm{d}t}{a}\,.$ (1)
The difference between the time $t$ of the observer and the time $\tau$ of the vacuum is the key to the Gibbons–Hawking effect [27]. For simplicity, we treat the electromagnetic field as a conformally coupled scalar field [27] $\widehat{A}$ with the correlation function
$K=\frac{1}{2}\langle 0|\widehat{A}_{1}\widehat{A}_{2}+\widehat{A}_{2}\widehat{A}_{1}|0\rangle$ (2)
of vacuum noise at the two space–time points $\{t_{1},{\bm{r}}_{1}\}$ and $\{t_{2},{\bm{r}}_{2}\}$ in co–moving coordinates indicated by the indices in $\widehat{A}_{1}$ and $\widehat{A}_{2}$. The vacuum energy density and pressure of the field can be calculated from $K$ by differentiating with respect to the space–time coordinates. Note that such a correlation function has been measured inside a dielectric material [32]. The correlation $K$ is the real part of the analytic function $\langle 0|\widehat{A}_{1}\widehat{A}_{2}|0\rangle$ while its imaginary part corresponds to the dissipation [31], the difference between the retarded and the advanced Green functions $G_{\pm}=-(4\pi r)^{-1}\delta(\tau_{21}\mp r/c)$ with $\tau_{21}=\tau_{2}-\tau_{1}$ and $r=|{\bm{r}}_{2}-{\bm{r}}_{1}|$. The fluctuation–dissipation theorem [31] relates the two, turning each delta function in the dissipation to a pole in the correlation. One obtains [23]:
$K=-\frac{1}{(2\pi)^{2}s^{2}}$ (3)
with
$s^{2}=a_{1}a_{2}\left[c^{2}(\tau_{2}-\tau_{1})^{2}-({\bm{r}}_{2}-{\bm{r}}_{1})^{2}\right].$ (4)
We treat the observer as a point detector at rest, ${\bm{r}}_{1}={\bm{r}}_{2}=\mathrm{const}$, sampling the correlation function $K$ along the proper time of the observer, the cosmological time $t$. We thus obtain from Eqs. (1), (3) and (4):
$K=-\frac{1}{(2\pi)^{2}c^{2}a_{1}a_{2}(\tau_{2}-\tau_{1})^{2}}=-\frac{1}{(2\pi)^{2}c^{2}}\,\frac{\dot{\tau}_{2}\,\dot{\tau}_{1}}{(\tau_{2}-\tau_{1})^{2}}$ (5)
with the dots denoting differentiation with respect to cosmological time. If we replace $\tau$ by the Möbius transformation
$\tau\rightarrow\frac{\alpha\,\tau+\beta}{\gamma\,\tau+\delta}$ (6)
with constant coefficients $\alpha$, $\beta$, $\gamma$, $\delta$ the correlation function (5) remains unchanged.
As an example, take the standard case of exponentially expanding space, de–Sitter space. There we have $a=a_{0}\mathrm{e}^{Ht}$ with Hubble constant $H$ and get from Eq. (1) the conformal time $\tau=-(aH)^{-1}$. This gives for $K$ formula (3) with
$s^{2}=\frac{c^{2}}{H^{2}}\,4\sinh^{2}\frac{H}{2}\theta\quad\mbox{and}\quad\theta=t_{2}-t_{1}\,.$ (7)
Consider now the thermal state in Minkowski space ($a=1$) with temperature $T$ written as
$k_{\mathrm{B}}T=\frac{\hbar H}{2\pi}\,.$ (8)
One gets from the fluctuation–dissipation theorem with Kubo–Martin–Schwinger relation [31] the correlation function:
$K_{\mathrm{th}}=-\frac{1}{8\pi^{2}c^{2}\rho^{2}}\,\partial_{\rho}\ln\left(\mathrm{e}^{H\theta}-\mathrm{a}^{H\rho}\right)\left(\mathrm{e}^{H\theta}-\mathrm{a}^{-H\rho}\right)\quad\mbox{with}\quad\rho=\frac{|{\bm{r}}_{2}-{\bm{r}}_{1}|}{c}\,.$ (9)
One verifies that $K$ in de–Sitter space [Eqs. (3) and (7)] agrees with $K_{\mathrm{th}}$ in Minkowski space for ${\bm{r}}_{1}={\bm{r}}_{2}$. This shows that the observer at rest in de–Sitter space perceives the quantum vacuum as a thermal state with Gibbons–Hawking temperature (8). Note that it has been essential here to treat the observer as a point detector, as the correlation function $K$ differs from $K_{\mathrm{th}}$ for ${\bm{r}}_{1}\neq{\bm{r}}_{2}$. Nonlocal detectors will respond differently to cosmic expansion. For example, one may take the ionisation of an atom by the expansion of space as a detection mechanism for Gibbons–Hawking radiation [33]. As the released electron tunnels out of the atom to infinity, this is not a local detection. For exponential expansion one gets exactly twice [33] the Gibbons–Hawking temperature of Eq. (8) — presumably because the expansion does not only facilitate but also accelerate the escape of the electron.
Finally, we turn from the Gibbons–Hawking effect of the observer at rest with the expanding universe to the Unruh effect of the accelerated observer. In the Unruh effect [4, 5, 6] a detector moves in empty Minkowski space on the space–time trajectory $t(\tau)$ and ${\bm{r}}(\tau)$ parameterized by the proper time of the detector, $\tau=\int\mathrm{d}\tau$ with $\mathrm{d}\tau^{2}=\mathrm{d}t^{2}-\mathrm{d}{\bm{r}}^{2}/c^{2}$. Here the roles of $t$ and $\tau$ are reversed: the vacuum propagates with Minkowski time $t$ while the spectrometer of the detector responds to the proper time $\tau$. The detector thus samples the correlation function (3) with
$s^{2}=\left[c^{2}(t_{2}-t_{1})^{2}-({\bm{r}}_{2}-{\bm{r}}_{1})^{2}\right].$ (10)
[…]
3 Wigner function
The spectrometer of the observer does two things: it decomposes the amplitude into spectral components and it detects them. As the spectrometer is attached to the observer, the spectrum is taken with respect to the proper time of the observer. This, in our opinion, is the most important aspect of the Gibbons–Hawking and the Unruh effect, not the actual act of detection [24]. It does not matter what exactly the detector does — it could measure the amplitude or count radiation quanta — as long as it is a point–detector. The spectrum is assumed to be continuously measured during the cosmic expansion or during the accelerated motion of the observer. For exponential expansion or uniform acceleration the correlation function depends only on the time difference. We could therefore describe the power spectrum by the Fourier transform with respect to that time difference. But in general, the correlation function will depend on two independent times. We could still define a time–dependent spectrum by Fourier–transforming with respect to the time difference:
$W=\frac{1}{2\pi}\int_{-\infty}^{+\infty}\mathrm{e}^{\mathrm{i}\omega\theta}K(t+\theta/2,t-\theta/2)\,\mathrm{d}\theta\,.$ (11)
This is the chronocyclic Wigner function adapted from classical [34] to quantum fields, the central subject of this paper. It was introduced before [23, 27], but in one study [23] it would produce highly oscillatory spectra when applied to cosmology and in the other [27] it was applied incorrectly for working out the high–frequency asymptotics of the spectrum. Here we will argue that it is a useful tool for low frequencies.
The Wigner function [28, 29, 30] was originally introduced [28] as a quasiprobability distribution for the position and the momentum of a quantum particle (and first applied in thermodynamics [28]). It can easily be generalized to other Fourier conjugates such as frequency and time, it has been generalized to angular–momentum variables [35] and to discrete variables [36, 37], and it is routinely measured in quantum state tomography [38, 39, 40]. For canonically conjugate variables, such as frequency and time, no proper phase–space distribution can exist, but the Wigner function acts as a good compromise. Its marginal distributions do give the correct power spectra or intensity distributions under linear transformations of the conjugate variables. This is its defining property [40].
In our case, the correlation function $K$ given by Eq. (5) is an even, real function with respect to the time difference $\theta=t_{2}-t_{1}$. From this follows that the Wigner function (11) is real and identical for ±$\pm\omega$. The same is true for the Wigner function of the accelerated observer. For an exponentially expanding universe (de–Sitter space) we get [23]:
$W=\frac{1}{(2\pi)^{2}c^{2}}\left(\frac{\omega}{2}+\frac{\omega}{\mathrm{e}^{2\pi\omega/H}-1}\right)\quad\mbox{for}\quad\omega>0$ (12)
and the same expression with $\omega$ replaced by $|\omega|$ for $\omega<0$. We may interpret this Wigner function as being proportional to the thermal energy of a bosonic point particle: the $\omega/2$ term describes the zero–point energy and the other term the Planck distribution of a thermal state with temperature (8). The point detector acts as a point particle in thermal equilibrium with Gibbons–Hawking temperature (8). Mathematically, the zero–point term appears due to the double pole of the correlation function (5) at $\theta=0$. We may remove this term by moving the integration contour in formula (11) infinitesimally above the real axis (Fig. 2) for $\omega>0$ (or below it for $\omega<0$).
This procedure has another advantage [27]. At the beginning of time, say at $t=0$, the scale factor develops a branch point [41] for a realistic universe (such that $a$ is no longer real or positive for $t<0$ — there was no discernible universe before the beginning of time. Suppose that $a\propto t^{1/\gamma}$ with constant $\gamma\geq 1$ for $t\sim 0$. From this follows [Eq. (5)] that $K\propto(t+\theta/2)^{-1/\gamma}+(t-\theta/2)^{-1/\gamma}$ as leading contribution near $\theta=\pm 2t$. We get from Wigner’s formula (11):
$W\propto\frac{\cos\left(2t\omega-\frac{\pi}{2\gamma}\right)}{\omega^{1-1/\gamma}}\quad\mbox{for}\quad\gamma\geq 1\,.$ (13)
Fourier transformations generate oscillations near sharp features; for the expanding universe we get spectral oscillations from the branch point of $a(t)$ at the beginning of time. Moving the integration contour slightly above or below the real axis avoids the branch points (Fig. 2) and thus removes this artefact. In a previous study [23] the integration was left on the real axis, which would have produced highly oscillatory, meaningless spectra if they were calculated.
Suppose, hypothetically, the universe were filled with one ideal fluid of energy density $\varepsilon$ and pressure $p$ with the equation of state $p=w\varepsilon$ and $w=\mathrm{const}$. From thermodynamics follows [41] $\mathrm{d}\varepsilon/\mathrm{d}a=-3(\varepsilon+p)/a=-3(1+w)\varepsilon/a$. Solving this differential equation gives $\varepsilon\propto a^{-2\gamma}$ with
$\gamma=\frac{3}{2}\,(1+w)\,.$ (14)
From the Friedmann equation [41] follows that $a\propto t^{1/\gamma}$. This is the case considered above. The amended Wigner function would be exactly zero, which is consistent with summing up all the phases in a cascade of Gibbons–Hawking processes at cosmological horizons [27]. They also give net zero radiation for a universe filled with a single fluid.
Consider now the realistic case of a fluid plus the cosmological constant (often called “the vacuum”). In this case the Friedmann equation [41] reads
$H^{2}=H_{\infty}^{2}\left(1+a^{-2\gamma}\right)$ (15)
with $\gamma$ given by Eq. (14) and $H$ being the Hubble parameter
$H=\frac{\dot{a}}{a}\,.$ (16)
The constant $H_{\infty}$ amounts to the Hubble parameter infinitely in the future when $a\rightarrow\infty$. For $w=0$ and hence $\gamma=3/2$ this model describes the current era of the universe turning from matter to vacuum domination. We obtain from the definition (16) of the Hubble parameter:
$t=\int\frac{\mathrm{d}a}{aH}\quad\implies\quad a=\left(\sinh\gamma H_{\infty}t\right)^{1/\gamma}\,.$ (17)
In the special case of $\gamma=1/2$ the scale factor turns into the square of the $\sinh$ of $H_{\infty}t/2$ and so the conformal time [Eq. (1)] becomes $\tau=-2H_{\infty}^{-1}\coth(H_{\infty}t/2)$. This is a Möbius transformation (6) of the conformal time for exponential expansion, $\tau=-H_{\infty}^{-1}\mathrm{e}^{-H_{\infty}t}$. We thus conclude that the Wigner function gives a perfect Planck spectrum for $\gamma=1/2$. For general $\gamma$ we obtain from the definition (1) of conformal time
$\tau=\int\frac{\mathrm{d}a}{a^{2}H}=H_{\infty}^{-1}\,\frac{a^{\gamma-1}}{\gamma-1}\,{}_{2}F_{1}\left(\frac{1}{2},\frac{1}{2}-\frac{1}{2\gamma},\frac{3}{2}-\frac{1}{2\gamma},-a^{2\gamma}\right)$ (18)
in terms of Gauss’ hypergeometric function ${}_{2}F_{1}$. We calculated numerically the Wigner function with the amended integration contour (Fig. 2) for a range of $\gamma$ parameters (with $\gamma>1$) and times $t$, and found excellent agreement with Planck spectra for low frequencies.
Figure 3 shows the results for the physically most relevant case, $\gamma=3/2$ (the current era of transition from matter to vacuum domination). The numerical data points agree with a Planckian fit well into the exponential tail of the spectrum (Fig. 3a). We use as fit
$W=\frac{A}{(2\pi)^{2}c^{2}}\,\frac{\omega}{\mathrm{e}^{2\pi\omega/H_{\mathrm{eff}}}-1}\,.$ (19)
characterized by the effective Hubble parameter $H_{\mathrm{eff}}$ for $H$ in the Gibbons–Hawking temperature (8) and the prefactor $A$. The effective temperature depends on time, but it does not vary much. If the spectra would vary rapidly in time, one could question whether they do represent actual spectra. Let us estimate this. The characteristic time $\Delta t$ over which the spectrum (19) varies is given by $1/\Delta t\sim|\partial_{t}H_{\mathrm{eff}}|/H_{\mathrm{eff}}$. The characteristic variation $\Delta\omega$ of the frequency in the Planck spectrum (19) is $\Delta\omega\sim H_{\mathrm{eff}}$. The spectrum would be affected by frequency–time uncertainty if $\Delta t\,\Delta\omega\sim 1$. However, our numerical results show that $|\partial H_{\mathrm{eff}}^{-1}/\partial t|\ll 1$ by about $10^{-2}$ and so $\Delta t\,\Delta\omega\gg 1$. The Planck spectra captured by the Wigner function (Fig. 3) are genuine for low frequencies.
For high frequencies, the spectrum becomes exponentially decaying and oscillatory. This is a remnant of the oscillations due to the branch point at the beginning of physical time, Eq. (13). By moving the integration contour away from the real axis we have eliminated this feature in the main part of the spectrum, but it reappears in the exponential tail. Figure 4 shows why. The integration contour can be deformed to go around the singularities and the branch cuts. The latter are situated on the lines $\mathrm{Re}\,\theta=\pm 2t$, $\mathrm{Im}\,\theta\geq 2\pi/\gamma$. The singularities produce, via Cauchy’s theorem, a geometrical series that sums up to a Planck spectrum with $H=H_{\infty}$:
$W_{\infty}=\frac{\omega}{(2\pi)^{2}c^{2}}\sum_{m=1}^{\infty}\mathrm{e}^{-2\pi\omega/H_{\infty}}=\frac{1}{(2\pi)^{2}c^{2}}\,\frac{\omega}{\mathrm{e}^{2\pi\omega/H_{\infty}}-1}\,.$ (20)
The branch cuts generate the corrections $W_{\mathrm{cuts}}$ to $W_{\infty}$ (as $W=W_{\infty}+W_{\mathrm{cuts}}$). For high frequencies and complex $\theta$ the Fourier factor $\mathrm{e}^{\mathrm{i}\omega\theta}$ in the Wigner integral (11) becomes exponentially decaying. For $\gamma>1$ the lowest branch point lies lower than the lowest singularity and so dominates the spectrum. Its imaginary part, $2\pi/\gamma$, gives an exponential decay, and its real part, $2t$, an oscillation. These features appear to be artefacts indicating the limitations of our Wigner method, but they only affect the exponential tail of the spectrum (Fig. 3).
Let us give another example showing a peculiar feature of the Wigner function (Fig. 5). Consider an analogue of cosmic expansion where the scale factor $a$ turns from a near–constant $a_{0}$ to another near–constant $a_{1}$ (Fig. 5a). For example, suppose
$a=\frac{a_{0}\,\mathrm{e}^{-H_{0}t}+a_{1}\,\mathrm{e}^{H_{0}t}}{\mathrm{e}^{-H_{0}t}+\mathrm{e}^{H_{0}t}}$ (21)
with $H_{0}=\mathrm{const}$. In this case, we obtain for the conformal time (1):
$\tau=\frac{(a_{0}+a_{1})H_{0}t+(a_{0}-a_{1})\ln(a_{0}\,\mathrm{e}^{-H_{0}t}+a_{1}\,\mathrm{e}^{H_{0}t})}{2a_{0}a_{1}H_{0}}\,.$ (22)
For $H_{0}t\ll-1$ the conformal time approaches $t/a_{0}$ and for $H_{0}t\gg+1$ it tends to $t/a_{1}$, both as it should for constant scale factors $a$ according to definition (1). Figure 5b shows the Wigner function. For $H_{0}t\lesssim-1$ we get a good Planck spectrum, but for $H_{0}t\gtrsim+1$ the Wigner function becomes negative. In quantum optics [30], negative regions of the Wigner function indicate nonclassical features [40]; we are not sure what they mean here.
4 Conclusions
We applied an amended [27] Wigner function [28, 29, 30] for calculating the spectrum of vacuum fluctuations for an observer at rest in the expanding universe or, equivalently, for an accelerated observer in a space at rest (in empty Minkowski space). The literature mostly focuses on exponential expansion (de–Sitter space) and uniform acceleration (Rindler trajectories) because these are the ideal cases where the cosmological or the Rindler horizons are genuine event horizons. Here we have developed tools for addressing the real cases of non–exponential expansion and non–uniform acceleration. The key feature that allows the calculation of physically meaningful spectra is the modification of the integration contour in the Wigner function [27]. Otherwise the spectra would be highly oscillatory and meaningless. For a class of models that includes the model describing the current cosmological era, we found excellent fits with Planck spectra. The vacuum noise in the expanding universe appears as thermal noise. However, we were not able to derive a simple analytic formula for the radiation temperature; a previous formula [27] for the high–frequency limit turned out to be incorrect. For high frequencies in the exponential tail of the spectrum, the Wigner function appears to exhibit spurious oscillations, and the Wigner method seems to fail there. For another type of cosmic evolution that can be experimentally tested in laboratory analogues [14, 15, 16, 17] we found non–Planckian, negative Wigner functions. In quantum optics [30] negative Wigner functions indicate non–classical features. What are they in our case? When is the spectrum Planckian? What is the temperature? When science answers one question, many more questions arise. We leave it to future readers to find some of the answers.
Acknowledgements
U.L. met Renaud Parentani on Capri in 2000, at a meeting on Cosmology in the Laboratory. While sitting together in the shade or strolling along the many paths on this beautiful island, talking physics, he learned from Renaud the beginning of the story that still occupies him today, the quantum physics of the cosmos. Many more meetings and discussions with Renaud were to follow, until a final one in Paris. His clarity and wit are missed now. We thank him for the gift he gave, and we also thank Michael Berry, Mathias Fink, Uwe Fischer, Amaury Micheli, Scott Robertson, Grisha Volovik and Chris Westbrook for discussions that contributed to this paper. The paper was supported by the Murray B. Koffler Professorial Chair at the Weizmann Institute.